# Probabilistic Feature Augmentation for AIS-Based Multi-Path Long-Term Vessel Trajectory Forecasting

### v5

@@ Abstract
Maritime transportation is paramount in achieving global economic growth, entailing concurrent ecological obligations in sustainability and safeguarding endangered marine species, most notably preserving large whale populations. In this regard, Automatic Identification System (AIS) data plays a significant role by offering real-time streaming data on vessel movement, allowing enhanced traffic monitoring. The impetus for the presented research was its use in a vessel trajectory forecast model module of a Decision Support System (DSS) that forecasts the risk of vessel-whale collisions. In this regard, this paper tackles an intrinsic problem to trajectory forecasting: the effective multi-path long-term vessel trajectory forecasting on engineered sequences of AIS data. For such a task, we have developed an encoder-decoder model architecture using Bidirectional Long Short-Term Memory Networks (Bi-LSTM) to predict the next 12 hours of vessel trajectories using 1 to 3 hours of AIS data as input. We feed the model with probabilistic features engineered from historical AIS data that refer to each trajectory’s potential route and destination. The model then predicts the vessel’s trajectory, considering these additional features by leveraging convolutional layers for spatial feature learning and a position-aware attention mechanism that increases the importance of recent timesteps of a sequence during temporal feature learning. The probabilistic features have an F1-Score of approximately 85% and 75% for each feature type, respectively, demonstrating their effectiveness in augmenting information to the neural network. While an earlier DSS version of the model already demonstrated strong predictive accuracy, the present version achieved an R2 score of over 98% in different techniques using varying features. The high R2 score is a natural outcome of the well-defined shipping lanes. However, our model stands out among other forecasting approaches as it demonstrates the capability of complex decision-making during turnings and path selection. Also, we have shown that our model can achieve more accurate forecasting with average and median forecasting errors of 11km and 6km, respectively. Our study demonstrates the potential use of data engineering and trajectory forecasting in marine species conservation efforts.
Introduction
@@
The International Maritime Organization (IMO) is essential in enhancing maritime safety, efficient navigation, and the prevention of marine pollution by ships [1]. One of the key technologies used to achieve these goals is the Automatic Identification System (AIS), which tracks marine vessels’ locations worldwide. Canadian regulations require every vessel planning to enter Canada to equip themselves with an AIS transceiver for sharing and receiving navigational data (Section 65 of the Navigation Safety Regulations111 Available at: https://laws-lois.justice.gc.ca/eng/regulations/SOR-2020-216). The AIS system uses a network of satellites and ground-based stations to track vessels’ location, speed, and other relevant data in real time. AIS allows for the real-time tracking of vessels, enhancing safety measures by improving the visibility of vessel movements for individual ships and shore-based facilities like the Coast Guard. By actively using AIS in maritime applications, authorities can track vessels, monitor their movements, and ensure compliance with regulations. These organizations can act in distress situations to prevent incidents, such as collisions [2, 3, 4], enhancing safety regardless of their current route (see Figure 1).
Research on AIS has opened the doors to enhance maritime safety and security through solving various issues such as the detection of suspicious events [5], monitoring traffic to avoid collisions [6, 7], and evaluating noise or environmental pollution levels [8, 9]. Recently, the problem of forecasting long-term trajectories [10, 11, 12, 13] of vessels gained significant attention to reduce vessel accidents, detect anomalous trajectories, and improve path planning — especially for autonomous shipping [14, 15, 16, 17, 18, 19, 20]. The study of trajectory forecasting has a diverse background, where different techniques have been employed to create models that predict movement, either from vessels or different vehicles. One of the most basic models is the constant velocity model [21], derived from the AIS messages’ geographical coordinates and the vessel’s Speed over Ground (SOG) and Course over Ground (COG) at the last recorded timestep, assuming no change in speed and course for the forecasting duration. Other models surpass these assumptions and use the trajectory’s geometry to produce more accurate predictions, such as the Ornstein–Uhlenbeck process adapted to vessel trajectory forecasting [22]. Machine and Deep Learning have advanced significantly in recent years, allowing for creating models that combine different types of data. However, the real-world use of these models is often limited due to the difficulty of synchronizing and preparing diverse data in streaming conditions.
@@
In a recent study by Nguyen et al. (2024) [23], authors proposed an approach to represent four AIS-related features into embeddings, i.e., a learned numerical representation of data where similar data points have a similar representation based on their meaning. Their study adopted the transformer encoder in a generative-driven approach to forecasting long-term trajectories, such as in Nguyen et al. (2021) [24]. However, their technique confines the ship’s trajectory to the centroid of the grid cells, requiring longer training sessions on a large model so it can learn the many patterns that distinguish each grid cell. This approach may be effective when ships adhere to nearly constant velocity models, but its potential to learn more individual moving patterns is overlooked. Besides this, the authors implemented their solution using teacher forcing [25]. Instead of inputting the previous step predictions as input to the next step, the technique trains the network with ground truth data, as that information is known during the training phase. During testing, the network is rewired to revert the need for ground truth data, expecting that the learned weights during forced training will perform better in the testing phase. To be more specific, the authors forecast information about the speed and course of the vessel, which can instead be mathematically calculated when having information about consecutive coordinates. Therefore, unless the model always provides a perfect estimation of the vessel’s speed and course, the information forwarded ahead in the model for predicting the next step of the sequence tends to be imprecise (i.e., uncertain).
@@
In different studies, Capobianco et al. (2021) [26] and Capobianco et al. (2022) [27] have reviewed the effectiveness of recurrent encoder-decoder neural networks in trajectory forecasting. In both studies, the authors utilized comparable network architectures that employ an encoder to produce an embedding; this embedding is then utilized in an attention mechanism to allow for multi-step forecasting and take advantage of both short and long-term dependencies learning, and the network decoder decodes the embedding and yields the forecasted trajectory. However, the issue in their modeling approach is not related to the auto-encoder model but rather to how the attention mechanism is employed. When attention is used over a sentence as seen in Large Language Models (LLMs) such as ChatGPT222 Available at: https://openai.com/chatgpt or Bert333 Available at: https://github.com/google-research/bert, the mechanism learns the significance (or importance) of words in the sentence, helping the model to focus on the key points of the sentence and respond accordingly to the LLM operator. However, sequences of AIS messages and sentences (i.e., sequences of words) have intrinsically different abstractions. In an AIS forecasting model, the knowledge that can be leveraged from the beginning of the sequence is considerably less important than the knowledge that can be extracted from later points in the same sequence because of the temporal and sequential nature of the data. This way, while the intermediate points of the sequence can aid in more accurate long-term trajectory forecasting, they also increase the uncertainty of the modeling problem in the short term because the model does not follow the correct order of events in the sequence.
@@
Using a different approach, this paper advances previous works by using an equal-size hexagon grid for extracting probabilistic features that aid in identifying the vessel route and destination. These features allow for a finer approximation of the vessel path without being constrained by the centroids of the cells. This means that differently from Nguyen et al. (2024) [23], we tackle this problem in a regression rather than a classification fashion, where the ocean grid is a step to achieve the trajectory instead of being an anchor for the trajectories themselves. Through this approach, we developed a model that can predict 12 hours of AIS data based on 3 hours of input data, where one AIS message is sent every 10 minutes. The study focuses on the Gulf of St. Lawrence in Canada, where vessels can take different routes to reach various regional ports. The ending point (destination) of each track is known, and the purpose is to identify the paths that cargo and tanker vessels will take in an area with many possible destinations (see Figure 1).
To this end, we propose a spatio-temporal model with an encoder-decoder structure where the encoder has three stacked blocks of convolutional layers attached to a Bidirectional Long Short-Term Memory Network (Bi-LSTM). Each block has a pair of $1\mathbf{D}$ convolutions where one extracts short-and the other longer-term dependencies from the input features. Before moving forward with the Bi-LSTM decoding, our model includes a positional-aware attention mechanism, which differs from Capobianco et al. (2021) [26] and Capobianco et al. (2022) [27] by allowing the auto-encoder model to learn from all AIS points in the trajectory but consistently giving more importance to the latter timestep of the input data sequence. The proposed modeling approach uses the attention mechanism as focal attention to the more recent AIS message for a smoother transition between inputs and outputs without requiring overly complex models.
As for the evaluation, we conducted an extensive set of experiments to test various scenarios and determine the effectiveness of our proposal. We employed different sets of features and applied trigonometric transformations to them, using the results to train models with varying complexities. In such a case, we observed that the effectiveness in providing probabilistic information to the neural network is in the order of 85% and 75% of the F1 Score in predicting the route and the destination of the vessels, respectively. Those results represent the degree of certainty of the probabilistic features. Furthermore, our findings indicate that when using only coordinates, speed, course information, and their deltas (see Methods) to predict long-term trajectories, the model’s ability to make path decisions is limited for cargo and tanker vessels. This is due to the model’s tendency to perform well on straight paths but struggle on complex and curved paths, resulting in mean and median errors of 13 km and 8 km, respectively, on the test dataset shared between both vessel types. Incorporating the proposed probabilistic features resulted in mean and median errors of 12 km and 8 km, respectively. However, after applying Cartesian and other trigonometrical transformations to all sets of features, the errors were reduced further, resulting in mean and median errors of 11 km and 6 km. By using the trigonometrical transformations on the probabilistic features, we were able to achieve an $R^{2}$ score higher than 98% for cargo and tankers using the proposed model, a notably high value that is due to the over occurrence of nearly straight paths in the dataset; even so, we demonstrate that our model stands out in decision-making in the face of curvy patterns and sudden route changes. Our contributions set a new milestone for long-term vessel trajectory forecasting using spatio-temporal data, probabilistic modeling of features, and advanced neural network techniques. These results have the potential to move forward with policies that can increase our awareness of the ocean, proactively aiding in reducing unintentional harm to endangered marine species, such as NARW.
@@
Figure 2: This study centers on the Gulf of St. Lawrence, with traffic policy zones defined by Transport Canada555 See: https://bit.ly/45DilrO.. These zones include voluntary seasonal slowdowns, dynamic shipping zones, and seasonal management areas, which are illustrated on the left-hand map. The regulated areas overlay main vessel traffic routes, such as the main shipping corridor between the Cabot Strait (between Cape Breton Island and Newfoundland) and the estuary of the St. Lawrence River routes (polygons 1–5). These areas cover the ports in the zone between New Brunswick and Québec (polygons 6–12) and the Jacques Cartier Strait between Anticosti Island and Northern Quebec (polygon 8). The visualization on the right-hand side displays vessel tracks (2015–2020) in the area, revealing inconsistent paths within shipping lanes.
@@
Problem Statement and Solution Formulation
The problem addressed in this paper and the achievements of the contributions were rooted in the Canadian Space Agency’s smartWhales program, which is a strategic response to the escalating issue of vessel-whale collisions in the Gulf of St. Lawrence. The related project (1 of 5 executed through the program), executed by WSP666 Available at: https://www.wsp.com/en-gl in collaboration with DHI Water & Environment777 Available at: https://www.dhigroup.com, Dalhousie University, the Institut des Sciences de la mer de Rimouski (ISMER)888 Available at: https://www.ismer.ca, and the Canadian Whale Institute999 Available at: https://www.canadianwhaleinstitute.ca, applies novel ecological modeling techniques and, among other data sources, Earth Observation Data to enable predictive hindcast and forecast modeling of NARW movements. This and the vessel trajectory forecast model technique described in this paper formed the foundation of a DSS capable of forecasting potential whale-life-threatening interactions between vessels and NARWs in the Gulf of St. Lawrence. The vessel-whale collisions issue, our underlying motivation, gained particular spotlight in 2017, a year marked by the reported fatalities of at least 12 NARW, highlighting the considerable risks posed by maritime activities, including shipping and fishing, to the survival of this species [29]. With the constant threat of collision and the persisting NARW endangered status, creating a model for forecasting multi-path long-term ship trajectories came as one of the ideas for aiding them in the minimization and solving of this issue.
With such a model in hand and real-time data acquisition, the smartWhales initiative can facilitate proactive mitigation steps that aim to minimize collision risks by providing advanced predictive analytics that anticipate the trajectories of vessels concerning the known habitats and movements of NARWs. In this sense, by adjusting vessel routes (see Figure 5B) avoid these critical areas (see Figure 5A), the smartWhales initiative aims to contribute to a significant reduction in the likelihood of harmful vessel-whale encounters mainly when vessel paths navigate across collision hotspot areas (see Figure 3). As vessels navigate, their historical data provide us with insights into their likely trajectory, along with observations on collective vessel movement patterns [7]; the same can be observed among people commuting [30, 31, 32], patient trajectories in hospitals [33], and spatio-temporal disease-spreading phenomena [34]. Accordingly, hereinafter, we describe how to leverage historical data in a probabilistic and machine-learning fashion to solve this problem.
@@
Accordingly, in the Gulf of St. Lawrence, ships can take various paths, even if their start and end are known beforehand (see Figure 1). This difficulty comes from the fact that no physical grid directs ships into specific lanes, unlike city streets [35, 36, 37]. Although traffic management policies designate shipping lanes, the lanes do not restrict the movement of vessels as they may enter or exit lanes at different points throughout the gulf. As a result, predicting ship trajectories becomes challenging, especially for longer distances, as the predicted paths deviate significantly from actual paths with increased elapsed time and travel distance. Therefore, we require a deep-learning model to learn from the input AIS data and its routing possibilities to understand the relationship between mobility constraints and moving patterns.
To solve this problem, we first split the gulf into hexagonal grid cells (see Figure 3) of $0.3^{\mathrm{o}}$ on EPSG:4269101010 See: https://epsg.io/4269. projection. The shape of the cells is due to the Earth’s curvature, as hexagonal cells can divide the Earth’s surface uniformly. Figure 4 shows some paths discovered from historical data, highlighted by polygons placed in high-traffic areas. Besides being placed in high-traffic areas, vessels crossing these polygons are more likely to interact with a whale hotspot area based on the historical data movement patterns in the same region. The intuition for utilizing grid cells is extracting the starting and destination area of tracks because the AIS data does not always provide detailed information regarding the destination of a voyage. Relationships between grid cells based on tracks will assist the decision-making of the AI model while forecasting trajectory in the absence of information related to the destination. In addition, there are countless possible locations for vessels in the gulf, making it difficult to compare them with others that have different, but yet similar, trajectories. Using grid cells helps us to overcome these challenges and allows for a more effective comparison between vessels that pass through the same grid cell. In this sense, comparing trajectories and the route polygons we placed in high-traffic areas, we can see that most tracks intersect with route polygons at some point during their course. However, paths like the one in Figure 4C are the most challenging ones in the dataset as it does not intersect with any route polygons and represent a curvy-shaped moving pattern. We developed a probabilistic model (see Algorithm 1) that leverages the relationship between grid cells, route polygons, and historical vessel tracks to predict the route a vessel will take and the destination (i.e., last grid cell) of the voyage. For the sake of disambiguation, this research proposes two models, one for engineering features and referred to as the probabilistic model, and another to forecast the trajectories, referred to as the deep-learning model.
@@
Understanding the historical movements and patterns of vessels is essential in this approach. Our probabilistic model, as shown in Figure 5B, considers the relationship of tracks with route polygons and grid cells that highlight major paths in the gulf. By calculating the (1) number of tracks going to a particular destination (a grid cell), (2) the direction of each track at the point where it intersects with a route polygon, and (3) the probability score of each surrounding cell as a potential destination for each track, we can gain valuable insights into ship behavior and the routes they take. The primary objective of this approach (i.e., probabilistic model) is to predict the probable destination (grid cell) while considering the vessel’s current position. This is achieved by computing the difference, through the Euclidean distance, between the motion behavior of the vessel in the grid cell where it lies and the historical motion behavior of vessels in each possible destination. We consider grid information and AIS messages to determine probability scores and potential destinations for a vessel. The output of this process is a set of engineered features that indicate the route and destination of a vessel for the problem of long-term multi-path trajectory forecasting.
Pipeline Overview
@@
In our trajectory forecasting pipeline, depicted in Figure 6, we extract MMSI, timestamps, vessel type, and coordinates from AIS messages recorded between 2015 and 2020. Although over ten types of vessels are included in the data, our study only focuses on cargo and tanker vessels due to their distinguishable linear movement behavior (Step 1). Our pipeline uses tracks, which are ordered sequences of a vessel’s positions from a source port to a destination, instead of complete voyages. Creating tracks from AIS messages involves segmenting voyages based on the time and distance between consecutive messages to identify sub-sequences with end-to-end trip behavior. In our study, a new track starts if the interval between two consecutive AIS messages exceeds 8 hours or the distance interval exceeds 50 Km. We weigh the tracks based on their starting and ending grid, assigning a weight value corresponding to the vessel density of that moving pattern with respect to the entire dataset. Such weight is used for stratified dataset sampling, assuring that distinct mobility patterns will be equally used for model training and validation (Step 2).

### v6

# Multi-Path Long-Term Vessel Trajectories Forecasting with Probabilistic Feature Fusion for Problem Shifting

@@ Abstract
This paper addresses the challenge of boosting the precision of multi-path long-term vessel trajectory forecasting on engineered sequences of Automatic Identification System (AIS) data using feature fusion for problem shifting. We have developed a deep auto-encoder model and a phased framework approach to predict the next 12 hours of vessel trajectories using 1 to 3 hours of AIS data as input. To this end, we fuse the spatiotemporal features from the AIS messages with probabilistic features engineered from historical AIS data referring to potential routes and destinations. As a result, we reduce the forecasting uncertainty by shifting the problem into a trajectory reconstruction problem. The probabilistic features have an F1-Score of approximately 85% and 75% for the vessel route and destination prediction, respectively. Under such circumstances, we achieved an R2 Score of over 98% with different layer structures and varying feature combinations; the high R2 Score is a natural outcome of the well-defined shipping lanes in the study region. However, our proposal stands out among competing approaches as it demonstrates the capability of complex decision-making during turnings and route selection. Furthermore, we have shown that our model achieves more accurate forecasting with average and median errors of 11km and 6km, respectively, a 25% improvement from the current state-of-the-art approaches. The resulting model from this proposal is deployed as part of a broader Decision Support System to safeguard whales by preventing the risk of vessel-whale collisions under the smartWhales initiative and acting on the Gulf of St. Lawrence in Atlantic Canada.
keywords:
Feature Fusion, Probabilistic Modeling, Deep Autoencoder, Spatiotemporal Forecasting, Phased Framework Approach, Trajectory Reconstruction
journal: Elsevier
url]www.spadon.com.br
[…]
1 Introduction & Background
The International Maritime Organization (IMO) is essential in enhancing maritime safety, efficient navigation, and the prevention of marine pollution by ships [1]. One of the key technologies used to achieve these goals is the Automatic Identification System (AIS), which tracks marine vessels’ locations worldwide. Canadian regulations require every vessel planning to enter Canada to equip themselves with an AIS transceiver for sharing and receiving navigational data (Section 65 of the Navigation Safety Regulations111 https://laws-lois.justice.gc.ca/eng/regulations/SOR-2020-216). The AIS system uses a network of satellites and ground-based stations to track vessels’ location, speed, and other relevant data in real-time. AIS allows for the real-time tracking of vessels, enhancing safety measures by improving the visibility of vessel movements for individual ships and shore-based facilities like the Coast Guard. By actively using AIS data in maritime applications, authorities can track vessels, monitor their movements, and ensure compliance with regulations. These organizations can act in distress situations to prevent incidents, such as collisions [2, 3, 4], enhancing safety regardless of their current route (see Figure 1).
Research on AIS data has opened the doors to enhance maritime safety and security through solving various issues such as the detection of suspicious events [5], monitoring traffic to avoid collisions [6, 7], and evaluating noise or environmental pollution levels [8, 9]. Recently, the problem of forecasting long-term trajectories [10, 11, 12, 13] of vessels gained significant attention to reduce vessel accidents, detect anomalous trajectories, and improve path planning — especially for autonomous shipping [14, 15, 16, 17, 18, 19, 20]. The study of trajectory forecasting has a diverse background, where different techniques have been employed to create models that predict vessel movement, either by observing a single vessel’s data or different vehicles. One of the most basic models is the constant velocity model [21], derived from the AIS messages’ geographical coordinates and the vessel’s Speed over Ground (SOG) and Course over Ground (COG) at the last recorded timestep, assuming no change in speed and course for the forecasting duration. Other models surpass these assumptions and use the trajectory’s geometry to produce predictions, such as the Ornstein–Uhlenbeck process is adapted for vessel trajectory forecasting [22]. Deep Learning has advanced significantly in recent years, allowing for the creation of models that combine different data types. However, the real-world use of these models is often limited due to the difficulty of synchronizing and preparing diverse data in streaming conditions.
@@ 1 Introduction & Background
In a recent study by Nguyen et al. (2024) [23], authors proposed an approach to represent four AIS-related features into embeddings, i.e., a learned numerical representation of data where similar data points have a similar representation based on their meaning. Their study adopted the transformer encoder in a generative-driven approach to forecasting long-term trajectories, as proposed by Nguyen et al. (2021) [24]. However, their technique confines the ship’s trajectory to the centroid of the grid cells, requiring longer training sessions on a large model so it can learn the many patterns that distinguish each grid cell. This approach may be effective when ships adhere to nearly constant velocity models, but its potential to learn more individual moving patterns is overlooked. Besides this, the authors implemented their solution using teacher forcing [25]. Instead of inputting the previous step predictions as input to the next step, the technique trains the network with ground truth data, as that information is known during the training phase. During testing, the network is rewired to revert the need for ground truth data, expecting that the learned weights during forced training will perform better in the testing phase. Specifically, the authors forecast information about the vessel’s speed and course, which can be mathematically calculated when having information about consecutive coordinates. Therefore, unless the model provides a perfect estimation of the vessel’s speed and course at every time step, the information used for predicting the next step of the sequence tends to be uncertain.
@@
In different studies, Capobianco et al. (2021) [26] and Capobianco et al. (2022) [27] have reviewed the effectiveness of recurrent encoder-decoder neural networks in trajectory forecasting. In both studies, the authors utilized comparable network architectures that employ an encoder to produce an embedding; this embedding is then utilized in an attention mechanism to allow for multi-step forecasting and take advantage of both short and long-term dependencies learning, and the network decoder decodes the embedding and yields the forecasted trajectory. However, the issue in their modeling approach is not related to the auto-encoder model but rather to how the attention mechanism is employed. When attention is used over a sentence as seen in Large Language Models (LLMs) such as ChatGPT222 https://openai.com/chatgpt or Bert333 https://github.com/google-research/bert, the mechanism learns the significance (or importance) of words in the sentence, helping the model to focus on the key points of the sentence and respond accordingly to the LLM operator. However, sequences of AIS messages and sentences (i.e., sequences of words) have intrinsically different abstractions. In an AIS forecasting model, the knowledge that can be leveraged from the beginning of the sequence is less important than the one that can be extracted from later points in the same sequence because of the spatiotemporal nature of the data. While the intermediate points of the sequence can aid in more accurate long-term trajectory forecasting, they also increase the uncertainty of the modeling problem in the short term because the model does not follow the correct order of events in the sequence.
@@
Using a different approach, this paper advances previous works by using an equal-size hexagon grid for extracting probabilistic features that aid in identifying the possible route and destination of different vessels — these features allow for a finer approximation of the vessel path without being constrained by the centroids of the cells. This means that, differently from Nguyen et al. (2024) [23], we tackle this problem in a regression rather than a classification fashion — where the ocean grid is used to distill spatial knowledge that will be fused with the existing trajectory data instead of working as an anchor for the trajectory forecasting task. Accordingly, our approach involves shifting the forecasting problem into a reconstruction by leveraging partial information about the future trajectory distilled from the grid system we previously created. Then, by fusing engineered data about the route and destination obtained from our spatial grid system with the AIS messages, we gain a deeper understanding of vessel behavior and can improve the precision of the forecasted (i.e., reconstructed) trajectory. As a key component of our approach, we developed a deep neural network model. This model has the capability to predict 12 hours of AIS data based on 3 hours of input data, with one AIS message being sent every 10 minutes. The model was designed to operate over the Gulf of St. Lawrence in Canada, where vessels can take different routes to reach various regional ports, with the purpose of identifying the paths that cargo and tanker vessels will take in an area with many possible destinations (see Figure 1).
The proposed model utilizes a spatio-temporal encoder-decoder architecture designed to enhance the trajectory reconstruction precision through phased feature learning and targeted attention mechanisms. The encoder comprises stacked CNN, each equipped with dual $1\mathbf{D}$ convolutions. These layers are specifically engineered to extract both short- and long-term dependencies from the input features using different kernel sizes, preparing a comprehensive spatial context for subsequent temporal processing. Following the CNN, the model employs a recurrent neural network (RNN) to handle the sequential data, which captures the temporal dynamics essential for understanding how the features evolve. This process is augmented with a positional-aware attention mechanism, pivotal in prioritizing more recent timesteps of the input sequence. Unlike the architectures described by Capobianco et al. (2021) [26] and Capobianco et al. (2022) [27], our model’s attention mechanism strategically emphasizes the latest data points, thereby ensuring they have a greater impact on the learning process. This methodological focus on the most current and relevant data points allows for a seamless and dynamic transition from input to output, enhancing the precision of the trajectory reconstruction task. By integrating these refined spatial and temporal features, our model produces accurate, timely, and contextually relevant predictions, reducing the uncertainty of long-term multi-path trajectory forecasting.
We conducted a comprehensive series of experiments to assess the performance of our model under various conditions. When assessing the capability of the spatial grid system to distill trustworthy information about the future location of a given trajectory, we achieved an F1-Score of approximately 85% and 75% for predicting vessel routes and destinations. At the same time, by excluding the engineered data from the spatial grid with the trajectory data and using basic features (such as coordinates, speed, and course information), the model’s decision-making capacity was limited. This was evident as the model excelled in predicting straightforward trajectories but struggled with complex or curved paths, initially showing mean and median errors of 13 km and 8 km, respectively. However, we significantly improved decision-making and reduced these errors by integrating probabilistic features showing mean and median errors of 11 km and 6 km, respectively. Remarkably, our solution achieved an $R^{2}$ score above 98% for cargo and tanker vessels. This high score primarily resulted from the frequent occurrence of nearly straight paths in the dataset. Nonetheless, our model demonstrated superior performance in navigating complex routes and responding to sudden changes in trajectory. Therefore, this research sets a new benchmark in long-term multi-path vessel trajectory forecasting and holds significant implications for marine safety and conservation efforts. By employing spatio-temporal data, probabilistic feature modeling, and sophisticated neural network techniques, we have paved the way for policies to increase marine safety and protect endangered species like the North Atlantic Right Whale (NARW) as a use case. The resulting model from this study is currently deployed as part of a broader Decision Support System to safeguard whales by preventing the risk of vessel-whale collisions under the smartWhales initiative and acting on the Gulf of St. Lawrence in Atlantic Canada.
This paper is structured into five sections besides the introduction, which offers an overview of the problem, existing literature, and the solutions we propose. Specifically, Section 2 covers the core problem of this research and outlines our solution that leverages a spatial grid system for knowledge extraction. Section 3 details the probabilistic model we developed to encapsulate spatial information pertinent to the routes and destinations of vessel trajectories. Section 4 discusses the deep neural network architecture tailored for trajectory reconstruction tasks. Section 5 presents our findings, providing a comprehensive discussion on the capacity of our proposed model. Finally, Section 6 summarizes our conclusions and final remarks. Supplementary concepts and definitions are provided in the Annex.
2 Problem Statement & Methodology
The problem addressed in this paper and the achievements of the contributions were rooted in the Canadian Space Agency’s smartWhales initiative, which is a strategic response to the escalating issue of vessel-whale collisions in the Gulf of St. Lawrence. The related project (1 of 5 executed through the program), executed by WSP444 https://www.wsp.com/en-gl in collaboration with DHI Water & Environment555 https://www.dhigroup.com, Dalhousie University, the Institut des Sciences de la mer de Rimouski (ISMER)666 https://www.ismer.ca, and the Canadian Whale Institute777 https://www.canadianwhaleinstitute.ca, applies novel ecological modeling techniques and, among other data sources, Earth Observation Data to enable predictive hindcast and forecast modeling of NARW movements. This and the vessel trajectory forecast model technique proposed in this paper formed the foundation of a Decision Support System capable of forecasting potential whale-life-threatening interactions between vessels and NARWs in the Gulf of St. Lawrence. The vessel-whale collisions issue, our underlying motivation, gained particular spotlight in 2017, a year marked by the reported fatalities of at least 12 NARW, highlighting the considerable risks posed by maritime activities, including shipping and fishing, to the survival of this species [28]. With the constant threat of collision and the persisting NARW endangered status, creating a model for forecasting multi-path long-term ship trajectories came as one of the ideas for aiding in solving this issue.
@@ 2 Problem Statement & Methodology
Figure 2: This study centers on the Gulf of St. Lawrence, with traffic policy zones defined by Transport Canada999 https://bit.ly/45DilrO. These zones include voluntary seasonal slowdowns, dynamic shipping zones, and seasonal management areas, which are illustrated on the left-hand map. The regulated areas overlay main vessel traffic routes, such as the main shipping corridor between the Cabot Strait (between Cape Breton Island and Newfoundland) and the estuary of the St. Lawrence River routes (polygons 1–5). These areas cover the ports in the zone between New Brunswick and Québec (polygons 6–12) and the Jacques Cartier Strait between Anticosti Island and Northern Quebec (polygon 8). The visualization on the right-hand side displays vessel tracks (2015–2020) in the area, revealing inconsistent paths within shipping lanes.
With such a model in hand and real-time data acquisition, the smartWhales initiative can facilitate proactive mitigation steps that aim to minimize collision risks by providing advanced predictive analytics that anticipate the trajectories of vessels concerning the known habitats and movements of NARWs. In this sense, by adjusting vessel routes (see Figure 9B) to avoid these critical areas (see Figure 9A), the smartWhales initiative aims to contribute to a significant reduction in the likelihood of harmful vessel-whale encounters mainly when vessel paths navigate across collision hotspot areas (see Figure 3). As vessels navigate, their historical data provide us with insights into their likely trajectory, along with observations on collective vessel movement patterns [7]; the same can be observed among people commuting [29, 30, 31], patient trajectories in hospitals [32], and disease-spreading phenomena [33].
@@
Accordingly, in the Gulf of St. Lawrence, ships can take various paths even if their start and end are known beforehand (see Figure 1). This is because no physical grid directs ships into specific lanes, unlike city streets [35, 36, 37]. Although traffic management policies designate shipping lanes, the lanes do not restrict the movement of vessels as they may enter or exit lanes at different points throughout the gulf. Predicting ship trajectories becomes increasingly challenging for longer distances as the predicted paths deviate significantly from actual paths with increased elapsed time and travel distance. In simpler terms, the further a ship travels, the larger the uncertainty of its future position, making it difficult to predict accurately. To overcome this challenge, we need a strategy that provides increased information for the forecasting task. This can be achieved by fusing information about a given vessel’s possible route and destination, which changes the forecasting paradigm into reconstruction while reducing the uncertainty of the vessel’s future positioning. However, such information is unavailable within the AIS data and would require a modeling strategy to provide trustworthy information about unseen and unknown trajectories in the ocean.
To solve this problem, we first split the gulf into hexagonal grid cells (see Figure 3) of $0.3^{\mathrm{o}}$ on EPSG:4269101010 https://epsg.io/4269. projection. The shape of the cells is due to the Earth’s curvature, as hexagonal cells can divide the Earth’s surface uniformly. On top of such grid, Figure 4 shows some paths discovered from historical data, highlighted by polygons placed in high-traffic areas. Vessels crossing these polygons are more likely to interact with a whale hotspot area based on the historical movement behaviors in the same region. In this sense, comparing trajectories in high-traffic areas, we can observe that most tracks intersect with route polygons at some point during their course. Besides, paths like the one in Figure 4C are the intricate patterns of the dataset as they do not intersect with any route polygons and represent curvy-shaped moving patterns. Based on all these different moving behaviors, we developed a probabilistic model (see Algorithm 1) that leverages the relationship between grid cells, route polygons, and historical vessel tracks to predict the route taken by a vessel and the destination (i.e., endpoint ) of that same voyage.
@@
Understanding the historical movements and patterns of vessels is essential in this approach. The proposed probabilistic model, as shown in Figure 5B, considers the relationship of tracks with route polygons and grid cells that highlight major paths in the gulf. By calculating the (1) number of tracks going to a particular destination (a grid cell), (2) the direction of each track at the point where it intersects with a route polygon, and (3) the probability score of each surrounding cell as a potential destination for each track, we can gain valuable insights into ship behavior and the routes they take. The initial step of this approach (i.e., probabilistic model) is to predict the probable destination (grid cell as endpoint) while considering the vessel’s current position. This is achieved by computing the difference, through the Euclidean distance (see Equation 1), between the motion behavior of the vessel in the grid cell where it currently lies and the historical motion behavior of vessels for each possible destination. We consider grid information and AIS messages to determine probability scores and potential destinations.
In the following, we provide a detailed description of the methodology we use to achieve the proposed solution, and for simplicity’s sake, we refer to forecasting and reconstruction indistinctly. Furthermore, for the sake of disambiguation, this research proposes two models: (1) a probabilistic model for predicting end-point of a track, and (2) deep-learning model to reconstruct the trajectories while using the output of probabilistic model as embedded features. Such models worked on a phased framework, contributing to the building of a feasible solution for the same problem.
Methodological Framework
@@
In our trajectory forecasting pipeline, as depicted in Figure 6, we extract MMSI, timestamps, vessel type, and coordinates from AIS messages recorded between 2015 and 2020. Although over ten types of vessels are included in the data, however, our study only focuses on cargo and tanker vessels due to their distinguishable quasi-linear movement behavior (Step 1). Our pipeline uses tracks, which are ordered sequences of a vessel’s positions from a source port to a destination, instead of complete voyages. Creating tracks from AIS messages involves segmenting voyages based on the time and distance between consecutive messages to identify sub-sequences with end-to-end trip behavior. In our study, a new track starts if the interval between two consecutive AIS messages exceeds 8 hours or the distance of 50 Km or change in the bearing between two consecutive points varies by more than ±$\pm 45^{g}$. It is obvious in the ocean that some routes are more busy than others thus each route has its own density of vessels. We weigh the tracks based on their starting and ending grid, assigning a weight value corresponding to the vessel density of that moving pattern with respect to the entire dataset. Such weight is used for stratified dataset sampling, assuring that distinct mobility patterns will be equally used for model training and validation (Step 2).